OpenStudy (anonymous):

simplify [25x^3/4] ^1/2 [49y^6]

OpenStudy (anonymous):

it's fraction?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

^ means to the power of

OpenStudy (anonymous):

\[\left(\frac{25x^{\frac{3}{4}}}{49y^6}\right)^\frac{1}{2}\] this?

OpenStudy (anonymous):

yeah, howd you type it like that?

OpenStudy (anonymous):

do you have an answer?

OpenStudy (mertsj):

5x^(3/8)/7y^3

OpenStudy (mertsj):